Fill in the boxes: 
  1 / 2 - ▯ = 1 / 6
Fill in the boxes: 
  1 / 2 - ▯ = 1 / 6

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

1 / 2 -▯= 1 / 6
▯=1 / 2–1 / 6
▯= [(1 × 3)–(1 × 1)] / 6
▯= (3–1) / 6
▯=2 / 6
▯ = 1 / 3

​Thanks